Systems Staff Meeting Notes 7/27/98

Systems Staff Meeting Notes
July 27, 1998

Present: Rich Bennett, Bill Covey, Debra Harris, Winston Harris, Sharon Henry, Bobbie Meng, Suzy Shaw, Leota Whitfield

Review Minutes of 7/20/98
  1. Bill (not Suzy) found the 'magic bullet fix': nbtstat -R
Announcements
  1. Xerox will return to the Libraries on August 20 for a group meeting - any problems we're experiencing should be presented at that time.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. MSL picked up 9 Cows. Further distribution will be: 28 new Cows (no zip drives) & 3 new Cows with zip drives. Also, 6 small machines to replace 3270 terminals.
  2. Iona came by & picked up a batch of Macs. John Humphrey indicated that the paperwork for the next batch is enroute to property, Eve needs to type it up, then Carolyn will sign - then it can go to PK Yonge. Once we get those out it will give us more space to take excess from various departments.
  3. Need to rectify Cow counts.
Liaisons
  1. Meeting next week.
  2. Need to reach a decision regarding Win'98 training. There is talk of upgrading Swamp II to Win'98, but upgrades are problematic. Microsoft denies that there is a systematic problem.
  3. Need to be on the lookout for the CIH virus that allegedly attacks the system at the BIOS level. Prom chips now have flash BIOS - some motherboards don't have it protected. This virus is capable of flashing the BIOS - supposedly on the 26th of each month. If the BIOS was attacked, a part would have to be ordered from the factory. It has all the earmarks of a hoax, but the possibility exists. Will have to come up with an upgrade to protect against the capabilities of such a virus.
  4. IBM 350's - memory upgrades not going so well - they don't seem to like EDO memory. Two have died after the memory installation - could have been done improperly, need to investigate.
Networking
  1. Public printing seems to be fixed. Bill had a meeting with Xerox this morning, their consultants cleaned up the left over mess from the previous installations & did a good job of it. There have been spotty reports of lost print jobs.
  2. The Novell Compaq server holds up the printing for the whole campus. The back up server is currently located on top of this one. It needs to be moved to the vault for safe-keeping. This server is plugged into it's own UPS & is not set for automatic power resurgence. If this building goes down, so does public printing.
  3. Walk-up access in room 100 - Capaldi has authorized $12-15,000. No wiring has been started yet. The networking plan for campus has not been set in that room yet.
  4. Special Collections - stack wiring is located at the wall. Wireless gizmo may be the answer. Bill will be flying in the Harris Corporation's private jet to Melbourne to look at their solution.
  5. Microsoft has not been very useful in solving our browser problem. The speed of sign-on was too slow - MS diagnosed it as a corrupted WINS database. Some parameters, which were static maps, were removed. Host map was installed on individual machines. The third silver bullet did not work, they are suppose to call back on Wednesday.
  6. LAC - their hub is port sensitive. Bill moved the cabling, may have to replace the hub or install switch. Hubs are not performing as they have before, especially the Boca hubs.
  7. POP server is not working on Exchange. Outbound mail isn't going out - it just goes to the Outbox. This isn't effecting everybody. The problem started over the weekend. Exchange upgrade may have to be done sooner than planned - this weekend or next.
  8. NERDC is retrenching ATM policy.
  9. Hacking is a serious matter - those caught will face jail time.
Software Projects Status
  1. Winston was able to connect to the FCLA database with ODBC. We could use MS Access or something similar. FCLA DB2 tables won't be stable for another month. Access may be via the micro rather than CICS. Will enable downloading to a server once a month. Potential cost is high. TSO charges may be reduced when VM goes away.
  2. Winston is splitting his time between SQL server Web access & ILL Borrowing.
  3. Evelyn seems to be doing well with the Personnel system - Sharon changed the authority to allow her to make more changes.
Calendar
  1. Winston out July 29 -31.
  2. Lee out Thursday, July 29.
  3. Debra out July 31, morning of August 3.
  4. Sharon out August 3.
  5. Lee out August 17-18.
  6. Bill out August 21 (Harris Corporation meeting).
  7. Lee out Sept. 17-30.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu